TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes


Created by Main.BorisBellesia on 2007-06-01 - 18:19

Powering Tests 7-8 - Evening Meeting 01/06/2007

ACTIVITY OF FRIDAY

* Triple quench in the matching section (Q5, Q4 and D2) is still under investigation. The current leads have been excluded as possible quench origin. New hardware will be implemented on Monday in order to better analyse the future events. In principle the quench in Q4 was not induced by the previous quench in D2, the delay is too short for that.

* The circuits in the DFBMH have been released by ElQA and QPS.

* 1.9K in the arc are not granted for Monday

PLAN FOR TODAY MONDAY

TEAM 1 - RR77 (& CCC)

* Connection of the four 600A current leads of the DFBMH (RQ6.R7B1 and RQ6.R7B2).

* PCC-Circuit Configuration tests in the circuits of DFBMH. Remark: the PC Permit signal from PIC will have to forced since the DFBMH belongs to the A78 powering subsector

* PIC2 in the DFBMH circuits (CCC)

TEAM 2 - CCC

* Powering of RQ5.L8 and its correctors to nominal current in intermidiate steps. Run for some hours at nominal current for validating the splices. If the test is successful the detection threshold of its quench detector will be risen to its normal operational value (100 mV)

ElQA

* If cryogenic conditions are recovered (2.1 K) ElQA of the RB.A78 circuit will continue. First, all the electronics of the DFBAs will be disconnected. If the test still fails the QPS electronics in the tunnel will be disconnected. Reiner ask for one hour advice for carrying out the disconnection.

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-06-01 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login